The length is 12.5 cm and the magnification is 70,000 X. calculate the length of this cell in µm = micro-meters.
Answer #1

Magnification = size of image/ actual size of object

70000=12.5cm/x

x=actual length of the cell  

x= 0.125 m/70000= 1.786*10^-6 m =1.786 micro-meters